### v1

@@ 5.1 Black Box Recorder evaluation
In evaluating the accountability component, we conducted six experiments across three previously described scenarios. These experiments involved integrity proofs storage in the blockchain, with hash calculations occurring at intervals of one every 10, 25, 50, or 100 messages. Moreover, our solution introduces the flexibility to save proofs of evidence at different rates based on the topic, ensuring non-tampering for critical topics or those with lower publishing rates where message loss could be crucial.
@@ 5.1.1 Metrics
Next, regarding system performance, we studied variables such as CPU, memory and disk usage. Specifically, we examined the following variables: system load 1-, 5- and 15-minute average, CPU usage (%), memory usage (GB), sent network traffic (MB/s) and disk writing rate (MB/s). These variables are highly significant for system-overhead analysis given that elevated load conditions can potentially influence the robot’s behavior and event-recording capabilities. In particular, we have obtained a full report about system load, CPU, RAM, network traffic and disk write operations updated every second, facilitating the tracking of overall system activity and an easy comparison with preceding values. To describe the overall distribution and characteristics of these results, we calculated their average as statistical measure. This information is included in Section 6.
@@ 5.2 Explainability evaluation
Despite the remarkable performance of LLMs across diverse language tasks, evaluating these models presents challenges, especially in ensuring factual consistency. The extensive capabilities of LLMs and the limitations of current benchmarks in capturing human preferences, have given rise to emerging trends in LLM evaluations, such as employing strong LLMs as judges (Zheng et al., 2023). Traditional similarity-based metrics, such as BLEU and ROUGE, have proven to misalign with human evaluation. While human evaluation remains the primary standard for assessing human preferences, it is exceptionally time-consuming and expensive. To automate the evaluation process, several works have adopted the ’LLM-as-a-judge’ approach, which focuses on using strong language models (LLMs) like GPT-4. These models are often trained with reinforcement learning from human feedback (RLHF), resulting in strong human alignment (Lin, 2024). Previous studies have evidenced the consistent agreement of these models with human grading in over 80% of judgments (Bubeck et al., 2023; Helgesson Hallström, 2023; Leng et al., 2023). Given these characteristics, we opted for GPT-4 to assess the answers generated by our explainability component. Through LLM-as-a-judge, we are able to obtain not only scores but also explanations, making their outputs interpretable. The whole process is represented in Figure 5.
@@
LLMs exhibit a certain degree of indeterminism, defined as the model’s ability to offer different answers when asked the same question or prompt several times. This behavior derived from their training, in which the models learn to generate outputs by sampling from probabilistic distributions over possible word sequences, introducing randomness to their response. The effect of this indeterminism must be considered in any LLM evaluation.
In addition, RAG for question-answering systems is one of the most difficult LLM use cases to evaluate. A basic RAG setup includes two main sources of error: the retriever and the LLM. The evaluation of the retriever aims at determining the relevance of its document selection in relation to a query. To this end, we conducted an experimental study running with 16 different questions over the data generated and processed during the navigation task in the three previously described scenarios. This set of questions has been evaluated six times in each scenario, giving a total of 288 questions and answers to perform the analysis. This approach ensures a thorough examination of the performance of the RAG system in various scenarios, showing potential strengths and weaknesses in our solution.
@@
A comprehensive evaluation of a LLM should consider multiple metrics that include correctness as well as other aspects. A language model should produce text that is not only correct but also logically structured and easy to understand. Produced response must be appropriately detailed without unnecessary verbosity. Furthermore, the LLM must not only provide correct information but also delivers in a meaningful and valuable way to the user.
Therefore, to assess the performance of the selected LLM on specific attributes, our approach also includes a criteria-based evaluation covering accuracy, conciseness, coherence, and helpfulness. Selected criteria metrics are describe in Table 3. This structured assessment offers a means to assess the model’s effectiveness in generating responses, providing a broader understanding of the model’s capabilities and its alignment with the used scenarios. Employing a scorecard defined for each of these features, our solution assigns scores between 0 and 10 to the responses it generates. The use of this scorecard facilitates a detailed and quantitative analysis of the LLM’s output, ensuring a comprehensive understanding of both its strengths and areas for improvement. A complete example of the answers and evaluations obtained for a selected question in the three selected scenarios is included in Figure 6.
@@
Criteria metric
Description
Accuracy
@@
Accuracy evaluates the extent to which the response conforms to the reference, ranging from being completely unrelated to being completely accurate and in line with the reference.
Helpfulness
@@
Helpfulness evaluates the degree to which the response provides valuable information, ranging from providing no helpful information to being extremely helpful and offering comprehensive information.
Conciseness
@@
Conciseness assesses the brevity of the response, ranging from being extremely verbose to highly concise with no unnecessary words.
Coherence
@@
Coherence examines the organizational structure and logical flow of the response, from extremely inconsistent to highly coherent, with a clear and logical organization.
@@
Question: How many goals have been reached by the robot? Scenario 1 Answer: Based on the provided context, we know that the robot has reached 3 goals. The context mentions that the goal number 1, 2 and 3 have been reached successfully. Therefore, the answer to your question is 3. GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The AI assistant’s response is accurate, concise, coherent, and helpful. It directly answers the user’s question by confirming that the robot has reached three goals. The assistant also provides clarification by referencing the context, which adds to the response’s accuracy and helpfulness. There are no unnecessary words or unclear statements, making the response concise and coherent. Accuracy: [[10]] Conciseness: [[10]] Coherence: [[10]] Helpfulness: [[10]]", score: 1.0 Scenario 2 Answer: Based on the context provided, the robot has reached 3 goals. GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The assistant’s response is completely accurate as it matches the information given in the ground truth. The response is concise, with no unnecessary words or information. It is highly coherent and has a clear structure. The response is extremely helpful as it gives the exact information asked for by the user. Therefore, the ratings are as follows: Accuracy: [[10]] Conciseness: [[10]] Coherence: [[10]] Helpfulness: [[10]]", score: 1.0 Scenario 3 Answer: Based on the provided context,the robot has successfully reached 2 out of 3 goals. The successful goals are: * Goal number 2: Navigation to the goal number 2 has succeeded. Position: -9.14068582195674, -25.977117025927537. Orientation: -0.6580310307304338,0.7529908117605705. * Goal number 3: Navigation to the goal number 3 has succeeded. Position: -1.8811226231679194, -28.773672979732826. Orientation: -0.5946646510401259,0.803973850820613. The aborted goal is: * Goal number 1: Navigation to the goal number 1 has been aborted. Position: -5.062160931312302, -8.330878461767831. Orientation: -0.8760794639388502,0.4821667479872674. Therefore, the answer to the user’s question is: The robot has successfully reached 2 out of 3 goals. Goal numbers 2 and 3 have been successfully achieved, while goal number 1 has been aborted. GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The AI assistant’s response is comprehensive and accurate. It correctly states that the robot has successfully reached goals number 2 and 3 while goal number 1 has been aborted. The assistant also provides additional details about the positions and orientations of the goals. While these extra details are not explicitly asked for in the user’s question, they may be useful for the user’s understanding. Therefore, the assistant’s response is also helpful. However, the response is not concise because it includes more information than necessary, leading to some redundancy in the explanation. The assistant’s response is coherent as it logically organizes the information about each goal. Accuracy: [[10]] Conciseness: [[7]] Coherence: [[10]] Helpfulness: [[10]]", score: 1.0
@@ 5.2.1 Categories and evaluation questions
To evaluate our explainability proposal in the context of ROS 2 messages interpretation, we have defined a set of questions covering diverse aspects of the navigation process over the how, the what, the which, the when and the where dimensions. These questions provide insights into different stages and components of the three scenarios. To structure and improve the understanding and analysis of this set of questions, we have organized them into five different categories. These categories range from high-level overviews of navigation to detailed inquiries about trajectory planning, goal completion, and the functioning of the behavior tree. The categories also delve into the causes of unexpected behavior or errors in navigation, improving the system explainability whenever an unexpected action occurs. Through these questions, we aim to address key concerns in XAI, including understanding why decisions are made, establishing causality between actions, determining success or failure, identifying timing, and disguising the methods by which achievements are realized (Gunning and Aha, 2019).
@@
Navigation Process Overview
@@
Trajectory Planning and Replanning
@@
    Did the robot find any obstacle during the navigation?
Goal Completion and Navigation Task Status
@@
Specifics about Goals and Locations
@@
    What was the linear velocity of the robot after receiving the goal number 1?
Nav2 Behavior Tree and Node Status
@@ 6.2 System benchmarking
Concerning memory consumption, all approaches within the three scenarios exhibit comparable values, differing by no more than 0.5 GB. Consequently, the inclusion of blockchain approaches into our black box recorder does not yield a remarkable impact on RAM comsuption in comparison to avoiding the use of anti-tampering techniques in this recording process.
Regarding network traffic, approaches with higher process storage in the blockchain increase this metric, but do not exceed 0,20 MB/s, which is a very affordable quantity. Furthermore, approaches such as storing one hash every 100 messages nearly match the avoidance of storing any proof into the blockchain, indicating that including this approach do not have a remarkable impact in the system performance.
Disk write operations do not form a bottleneck in system performance, with all approaches differing by no more than 0,40 MB/s, which is not a very significant impact. In addition, all the approaches that perform integrity proof storage on the blockchain achieve lower or similar results to the approach avoiding anti-tampering operations.
@@ 6.3 Correctness evaluation
Figure 10 shows the results of the correctness of RAG evaluation, presenting both the overall results and assessments based on predefined categories. The results demonstrate remarkable consistency across the three scenarios, getting a significant number of correct answers, even though there has not been a prior tuning in the ROS 2 messages for autonomous agents navigation tasks. The slight percentage of incorrect answers remains within an acceptable threshold, given the specific context employed and the inclusion of general questions.
@@
Regarding individual categories, the correctness in the category ’Navigation Process Overview’ is relatively consistent across all scenarios, indicating a good understanding of events and progress related to navigation. However, the broad nature of the question "What has happened in this ROS2 log regarding navigation?" initially yielded a correctness below 65% in the third scenario. In this instance, the cancellation of a goal resulted in the execution and failures of Nav2 Behavior Tree nodes, not included in the other two scenarios and focused on navigation recovery. Therefore, the incorporation in a dataset of information related to the cancellation of a goal, results in LLM asnwers that usually diverge from the ground truth provided, as the model occasionally prioritized Behavior Trees information, other times focused on locations or timestamps. This variability made it difficult to align with the ground truth provided, even if the obtained answers could be considered partially accurate. Introducing the prompt extension "Describe the sequence of events regarding navigation goals and include information about each of them." in question 1 significantly increased correctness in the first category, achieving results around and exceeding the 75% in all three scenarios. This fact underscores the effectiveness of prompt engineering when tackling ambiguous or broad queries.
@@
The results also show consistently high correctness across all scenarios concerning questions related to trajectory planning and re-planning. The model is able to determine whether the modification of a predicted trajectory has been caused by the occurrence of an obstacle or other circumstance. Scenario 1 and Scenario 2 also exhibit relatively high correctness in questions related to goal completion and task status. However, the model encounters challenges in Scenario 3, occasionally failing to identify if a navigation that starts correctly concludes with an abortion or cancellation. This issue may be attributed to misalignment during the retrieval process, specifically in the selection of the best chunk associated with a particular goal associated with numerous embeddings within a single collection from the vector database.
@@
Correctness in identifying specific targets and locations lacks some consistency, with results exceeding 50% in the first scenario, around 70% in the second and over 75% in the third. In particular, three scenarios encounter difficulties in identifying the initial position of the robot before starting the navigation, due to the model misaligns the queried position with the one assumed by the robot when starting the navigation process. This fact may also be attributed to deficiencies in the retrieval process. In Scenario 1, understanding the sequence of events or the order of goals results challenging for the model, leading to inaccuracies in specifying the second location or goal. Nevertheless, this challenge is not present in Scenario 3 and absent or reduced in Scenario 2, demonstrating how differences in data distribution can introduce varied responses even when using specific and unambiguous questions.
@@
Through questions included in the Nav2 Behavior Tree and node status category, we assessed the effectiveness of the model’s formal linguistic competence compared to its functional linguistic competence, a feature of present LLMs (Mahowald et al., 2023). While the model excels in identifying Behavior Tree Nodes given its description, it faces challenges in pinpointing these components under specific conditions. As a result, the category correctness suffers a drop in Scenario 3, although overall results remain above 75%. In this scenario, the navigation recovery actions caused by the cancellation of the first goal introduced a challenge in the interpretation of the information by the LLM due to the execution of several Nav2 Behavior Tree nodes not present in the others scenarios. Particularly, the mistakes stem from the model’s confusion in determining whether a previously failed Nav2 Behavior Tree node has recovered or finished its execution, as well as identifying all nodes that were executed during navigation. This can be attributed to the fact that the complexities of the recovery actions and the dynamic nature of Behavior Tree executions during navigation recovery introduce complexities that the model may not fully generalize from the training data, underscoring the need for additional fine-tuning in such scenario-specific contexts in case the obtained correctness needs to be improved.

### v2

@@ 5.1 Black Box Recorder Evaluation
In evaluating the accountability component, we conducted six experiments across three previously described scenarios. These experiments involved storing integrity proofs on the blockchain, with hash calculations occurring at intervals of one every 10, 25, 50, or 100 messages. Moreover, our solution introduces the flexibility to save proofs of evidence at different rates based on the topic, ensuring non-tampering for critical topics or those with lower publishing rates where message loss could be crucial.
@@ 5.1.1 Metrics
Next, regarding system performance, we studied variables such as CPU, memory and disk usage. Specifically, we examined the following variables: system load 1-, 5- and 15-minute average, CPU usage (%), memory usage (GB), sent network traffic (MB/s) and disk writing rate (MB/s). These variables are highly significant for system-overhead analysis given that elevated load conditions can potentially influence the robot’s behavior and event-recording capabilities. In particular, we have obtained a full report about system load, CPU, RAM, network traffic and disk write operations updated every second, facilitating the tracking of overall system activity and an easy comparison with preceding values. To describe the overall distribution and characteristics of these results, we calculated their average as a statistical measure. This information is included in Section 6.
@@ 5.2 Explainability Evaluation
Despite the remarkable performance of LLMs across diverse language tasks, evaluating these models presents challenges, especially in ensuring factual consistency. The extensive capabilities of LLMs and the limitations of current benchmarks in capturing human preferences have given rise to emerging trends in LLM evaluations, such as employing strong LLMs as judges [59]. Traditional similarity-based metrics, such as BLEU and ROUGE, have proven to misalign with human evaluation. While human evaluation remains the primary standard for assessing human preferences, it is exceptionally time-consuming and expensive. To automate the evaluation process, several works have adopted the ’LLM-as-a-judge’ approach, which focuses on using strong LLMs like GPT-4. These models are often trained with RLHF, resulting in strong human alignment [75]. Previous studies have evidenced the consistent agreement of these models with human grading in over 80% of judgments [76, 77, 78]. Given these characteristics, we opted for GPT-4 to assess the answers generated by our explainability component. Through LLM-as-a-judge, we can obtain not only scores but also explanations, making their outputs interpretable. The whole process is represented in Figure 5.
@@
LLMs exhibit a certain degree of indeterminism, defined as the model’s ability to offer different answers when asked the same question or prompt several times. This behavior derives from their training, in which the models learn to generate outputs by sampling from probabilistic distributions over possible word sequences, introducing randomness to their response. The effect of this indeterminism must be considered in any LLM evaluation.
In addition, RAG for question-answering systems is one of the most difficult LLM use cases to evaluate. A basic RAG setup includes two main sources of error: the retriever and the LLM. The evaluation of the retriever aims to determine the relevance of its document selection concerning a query. To this end, we conducted an experimental study with 16 different questions on the data generated and processed during the navigation task in the three previously described scenarios. This set of questions has been evaluated six times in each scenario, giving a total of 288 questions and answers to perform the analysis. This approach ensures a thorough examination of the performance of the RAG system in various scenarios, showing potential strengths and weaknesses in our solution.
@@
A comprehensive evaluation of an LLM should consider multiple metrics that include correctness as well as other aspects. A language model should produce text that is not only correct but also logically structured and easy to understand. Produced responses must be appropriately detailed without unnecessary verbosity. Furthermore, the LLM must not only provide correct information but also deliver it in a meaningful and valuable way to the user.
Therefore, to assess the performance of the selected LLM on specific attributes, our approach also includes a criteria-based evaluation covering accuracy, conciseness, coherence, and helpfulness. Selected criteria metrics are described in Table 3. This structured assessment offers a means to assess the model’s effectiveness in generating responses, providing a broader understanding of the model’s capabilities and its alignment with the used scenarios. Employing a scorecard defined for each of these features, our solution assigns scores between 0 and 10 to the responses it generates. The use of this scorecard facilitates a detailed and quantitative analysis of the LLM’s output, ensuring a comprehensive understanding of both its strengths and areas for improvement. A complete example of the answers and evaluations obtained for a selected question in the three selected scenarios is included in Figure 6.
@@
Criteria metric Description
@@
Accuracy Accuracy evaluates the extent to which the response conforms to the reference, ranging from being completely unrelated to being completely accurate and in line with the reference.
@@
Helpfulness Helpfulness evaluates the degree to which the response provides valuable information, ranging from providing no helpful information to being extremely helpful and offering comprehensive information.
@@
Conciseness Conciseness assesses the brevity of the response, ranging from being extremely verbose to highly concise with no unnecessary words.
@@
Coherence Coherence examines the organizational structure and logical flow of the response, from extremely inconsistent to highly coherent, with a clear and logical organization.
@@
{mdframed} [hidealllines=true,backgroundcolor=blue!20,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt, skipabove=-3pt,skipbelow=-10pt] Question: How many goals have been reached by the robot? {mdframed}[hidealllines=true,backgroundcolor=blue!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] Scenario 1 Answer: Based on the provided context, we know that the robot has reached 3 goals. The context mentions that the goal number 1, 2 and 3 have been reached successfully. Therefore, the answer to your question is 3.{mdframed} [hidealllines=true,backgroundcolor=gray!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The AI assistant’s response is accurate, concise, coherent, and helpful. It directly answers the user’s question by confirming that the robot has reached three goals. The assistant also provides clarification by referencing the context, which adds to the response’s accuracy and helpfulness. There are no unnecessary words or unclear statements, making the response concise and coherent. Accuracy: [[10]] Conciseness: [[10]] Coherence: [[10]] Helpfulness: [[10]]", score: 1.0{mdframed} [hidealllines=true,backgroundcolor=blue!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] Scenario 2 Answer: Based on the context provided, the robot has reached 3 goals.{mdframed} [hidealllines=true,backgroundcolor=gray!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The assistant’s response is completely accurate as it matches the information given in the ground truth. The response is concise, with no unnecessary words or information. It is highly coherent and has a clear structure. The response is extremely helpful as it gives the exact information asked for by the user. Therefore, the ratings are as follows: Accuracy: [[10]] Conciseness: [[10]] Coherence: [[10]] Helpfulness: [[10]]", score: 1.0{mdframed} [hidealllines=true,backgroundcolor=blue!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] Scenario 3 Answer: Based on the provided context,the robot has successfully reached 2 out of 3 goals. The successful goals are: * Goal number 2: Navigation to the goal number 2 has succeeded. Position: -9.14068582195674, -25.977117025927537. Orientation: -0.6580310307304338,0.7529908117605705. * Goal number 3: Navigation to the goal number 3 has succeeded. Position: -1.8811226231679194, -28.773672979732826. Orientation: -0.5946646510401259,0.803973850820613. The aborted goal is: * Goal number 1: Navigation to the goal number 1 has been aborted. Position: -5.062160931312302, -8.330878461767831. Orientation: -0.8760794639388502,0.4821667479872674. Therefore, the answer to the user’s question is: The robot has successfully reached 2 out of 3 goals. Goal numbers 2 and 3 have been successfully achieved, while goal number 1 has been aborted.{mdframed} [hidealllines=true,backgroundcolor=gray!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The AI assistant’s response is comprehensive and accurate. It correctly states that the robot has successfully reached goals number 2 and 3 while goal number 1 has been aborted. The assistant also provides additional details about the positions and orientations of the goals. While these extra details are not explicitly asked for in the user’s question, they may be useful for the user’s understanding. Therefore, the assistant’s response is also helpful. However, the response is not concise because it includes more information than necessary, leading to some redundancy in the explanation. The assistant’s response is coherent as it logically organizes the information about each goal. Accuracy: [[10]] Conciseness: [[7]] Coherence: [[10]] Helpfulness: [[10]]", score: 1.0
@@ 5.2.1 Categories and Evaluation Questions
To evaluate our explainability proposal in the context of ROS 2 messages interpretation, we have defined a set of questions covering diverse aspects of the navigation process over the how, the what, the which, the when and the where dimensions. These questions provide insights into the different stages and components of the three scenarios. To structure and improve the understanding and analysis of this set of questions, we have organized them into five different categories. These categories range from high-level overviews of navigation to detailed inquiries about trajectory planning, goal completion, and the functioning of the behavior tree. The categories also delve into the causes of unexpected behavior or errors in navigation, improving the explainability of the system whenever an unexpected action occurs. Through these questions, we aim to address key concerns in XAI, including understanding why decisions are made, establishing causality between actions, determining success or failure, identifying timing, and disguising the methods by which achievements are realized [79].
@@
Navigation Process Overview.
@@
Trajectory Planning and Re-planning.
@@
    Did the robot find any obstacles during the navigation?
Goal Completion and Navigation Task Status.
@@
Specifics about Goals and Locations.
@@
    What was the linear velocity of the robot after receiving goal number 1?
Nav2 Behavior Tree and Node Status.
@@ 6.2 System Benchmarking
Concerning memory consumption, all approaches within the three scenarios exhibit comparable values, differing by no more than 0.5 GB. Consequently, the inclusion of blockchain approaches into our black box recorder does not yield a remarkable impact on RAM consumption in comparison to avoiding the use of anti-tampering techniques in this recording process.
Regarding network traffic, approaches with higher process storage in the blockchain increase this metric, but do not exceed 0,20 MB/s, which is a very affordable quantity. Furthermore, approaches such as storing one hash every 100 messages nearly match the avoidance of storing any proof into the blockchain, indicating that including this approach does not have a remarkable impact on the system performance.
Disk write operations do not form a bottleneck in system performance, with all approaches differing by no more than 0,40 MB/s, which is not a very significant impact. In addition, all the approaches that perform integrity-proof storage on the blockchain achieve lower or similar results to the approach avoiding anti-tampering operations.
@@ 6.3 Correctness Evaluation
Figure 10 shows the results of the correctness of RAG evaluation, presenting both the overall results and assessments based on predefined categories. The results demonstrate remarkable consistency across the three scenarios, getting a significant number of correct answers, even though there has not been a prior tuning in the ROS 2 messages for autonomous agents’ navigation tasks. The slight percentage of incorrect answers remains within an acceptable threshold, given the specific context employed and the inclusion of general questions.
@@
Regarding individual categories, the correctness in the category ’Navigation Process Overview’ is relatively consistent across all scenarios, indicating a good understanding of events and progress related to navigation. However, the broad nature of the question "What has happened in this ROS 2 log regarding navigation?" initially yielded a correctness below 65% in the third scenario. In this instance, the cancellation of a goal resulted in the execution and failures of Nav2 Behavior Tree nodes, not included in the other two scenarios and focused on navigation recovery. Therefore, the incorporation in a dataset of information related to the cancellation of a goal results in LLM answers that usually diverge from the ground truth provided, as the model occasionally prioritized Behavior Trees information, other times focused on locations or timestamps. This variability made it difficult to align with the ground truth provided, even if the obtained answers could be considered partially accurate. Introducing the prompt extension "Describe the sequence of events regarding navigation goals and include information about each of them." in question 1 significantly increased correctness in the first category, achieving results around and exceeding 75% in all three scenarios. This fact underscores the effectiveness of prompt engineering when tackling ambiguous or broad queries.
@@
The results also show consistently high correctness across all scenarios concerning questions related to trajectory planning and re-planning. The model can determine whether the modification of a predicted trajectory has been caused by the occurrence of an obstacle or other circumstance. Scenario 1 and Scenario 2 also exhibit relatively high correctness in questions related to goal completion and task status. However, the model encounters challenges in Scenario 3, occasionally failing to identify whether a navigation that starts correctly concludes with an abortion or cancellation. This issue may be attributed to misalignment during the retrieval process, specifically in the selection of the best chunk associated with a particular goal associated with numerous embeddings within a single collection from the vector database.
@@
Correctness in identifying specific targets and locations lacks some consistency, with results exceeding 50% in the first scenario, around 70% in the second and over 75% in the third. In particular, three scenarios encounter difficulties in identifying the initial position of the robot before starting the navigation, as the model misaligns the queried position with the one assumed by the robot when starting the navigation process. This fact may also be attributed to deficiencies in the retrieval process. In Scenario 1, understanding the sequence of events or the order of goals is a challenge for the model, leading to inaccuracies in specifying the second location or goal. Nevertheless, this challenge is not present in Scenario 3 and is absent or reduced in Scenario 2, demonstrating how differences in data distribution can introduce varied responses even when using specific and unambiguous questions.
@@
Through questions included in the Nav2 Behavior Tree and node status category, we assessed the effectiveness of the model’s formal linguistic competence compared to its functional linguistic competence, a feature of present LLMs [80]. While the model excels in identifying Behavior Tree Nodes given its description, it faces challenges in pinpointing these components under specific conditions. As a result, the category correctness suffers a drop in Scenario 3, although overall results remain above 75%. In this scenario, the navigation recovery actions caused by the cancellation of the first goal introduced a challenge in the interpretation of the information by the LLM due to the execution of several Nav2 Behavior Tree nodes not present in the other scenarios. Particularly, the mistakes stem from the model’s confusion in determining whether a previously failed Nav2 Behavior Tree node has recovered or finished its execution, as well as identifying all nodes that were executed during navigation. This can be attributed to the fact that the complexities of the recovery actions and the dynamic nature of Behavior Tree executions during navigation recovery introduce complexities that the model may not fully generalize from the training data, underscoring the need for additional fine-tuning in such scenario-specific contexts in case the obtained correctness needs to be improved.